# LoRA vs Full Fine-tuning: An Illusion of Equivalence

### v2

@@ 3 Structural Differences
1:Pre-trained weights $W_{0}$, fine-tuned weights $W_{\text{t}}$, cosine similarity threshold $\epsilon$, # of fine-tuned singular vectors to examine $k$.
2:$[U_{0},\Sigma_{0},V_{0}^{\!\top}]\leftarrow\text{SVD}(W_{0})$
3:$[U_{t},\Sigma_{t},V_{t}^{\!\top}]\leftarrow\text{SVD}(W_{\text{tuned}})$
4:${{\rm n\_intruders}}\leftarrow 0$
5:$n\leftarrow$ of pre-trained singular vectors
@@
LoRA fine-tuned models contain high-ranking intruder dimensions while fully fine-tuned models do not. To characterize the differences in fine-tuning methods, we first evaluate the differences in the total number of intruder dimensions in the top 10 highest-ranking singular vectors ($k=10$). We repeat this procedure for a the range of $\epsilon$ values, our cosine similarity threshold. The results are presented in Fig. 4. For LLaMA2-7B, we find that models trained with LoRA contain intruder dimensions for ranks at least as high as $r\leq 256$. For RoBERTa, we consistently observe intruder dimensions for rank $r\leq 16$, even for low values of $\epsilon$. Interestingly, we observe that fully fine-tuned models, for all model sizes, almost never contain intruder dimensions in their top 10 singular vectors, even for epsilon values of about 0.6 to 0.9. This means that full fine-tuning makes smaller changes to the same set of high contribution pre-trained singular vectors, rather than introducing new singular vectors like LoRA. Importantly, the number of intruder dimensions appears to drop as rank increases past a certain threshold, suggesting that the low-rank nature, as well as the update rule of LoRA, induces them to occur. This is underscored by the $r=2048$ case of LLaMA2-7B fine-tuned on math (Fig 4(b)), which does not have intruder dimensions and instead has a very similar curve to full fine-tuning. As rank increases past a threshold and LoRA begins to resemble a high rank update, intruder dimensions begin to disappear.
LoRA variants have intruder dimensions. We examine 4 other LoRA variants to ensure that our findings to not just apply to vanilla LoRA. We examine AdaLoRA (Zhang et al., 2023), LoRA+ (Hayou et al., 2024), PiSSA (Meng et al., 2024), and VeRA (Kopiczko et al., 2024). In all of these cases, we find intruder dimensions with similar characteristics to vanilla LoRA (see Fig. 23). This shows our findings hold to other variants. For more discussion about these methods, see Appendix O.
@@ 4 Model Differences: Forgetting and Out-of-Distribution Generalization
LoRA forgets less. We measure the change in out of distribution performance (forgetting) induced by fine-tuning. For LLaMA2-7B, we follow Biderman et al. (2024) and measure forgetting as the average score on Hellaswag (Zellers et al., 2019), WinoGrande (Sakaguchi et al., 2021), Arc-Challenge (Clark et al., 2018). For RoBERTa-base, we measure its “pseudo-loss”, which is analogous to language modelling loss for encoder-only models, as described by Salazar et al. (2020) on a sample of its pre-training dataset (as described by Liu et al. (2019)). Going forward, we refer to these values as “forgetting” and report them in Fig. 6. We observe that across all tasks, full fine-tuning forgets more of its pre-training language modeling ability in comparison to LoRA. Importantly, all our RoBERTa-base models fine-tune to equivalent accuracy on the downstream task (Table 2). This extends the finding that LoRA learns less Biderman et al. (2024) to the case where LoRA and full fine-tuning have equal fit, showing that LoRA forgetting less is not simply a function of it underfitting the fine-tuning task in comparison to full fine-tuning (like in Biderman et al. (2024)), but rather a characteristic of LoRA itself.
@@
Intruder dimensions strongly correlate with forgetting. When we measure the Spearman correlation between the number of intruder dimensions with forgetting in Fig. 7, we find an extremely strong fit($\rho=0.971$, p-value $\ll 0.001$). When measuring the same for our LLaMA2-7B models across training epochs (Fig. 11), we still find a strong and still statistically significant relationship ($\rho=0.59$, p-value $=0.0006$). In contrast, when measuring the correlation between intruder dimensions and test accuracy, we find no statistically significant relationship: for RoBERTa, we measure $\rho=-0.3381$ p-value $=0.218$. For LLaMA2-7B, we measure $\rho=-0.3178$ p-value $=0.0869$. See Appendix F for more information. These results suggest that intruder dimensions are clearly linked with forgetting but are not necessary for performance. We examine this claim and whether this relationship is causal in the next section.
@@ A.2 Adding a random vector
Adding an random vector to a pre-trained matrix introduces an intruder dimension: To help provide intuition about how new singular vectors in the SVD can be added by LoRA, we examine mathematical conditions that lead to their creation. Certainly, when comparing SVD$(W+\lambda vv^{T})$ and SVD$(W)$, where $W$ are the pre-trained weights in $\mathbb{R}^{n\times n}$, $v$ is a randomly sampled vector in $\mathbb{R}^{n}$, and $\lambda$ is a scalar value greater than the largest singular value of $W$, we expect this update to create an intruder dimension (as $v$ is nearly orthogonal to the existing singular vectors w.h.p.).
@@ Appendix C Model Accuracies
RoBERTa-base Full 0.8617 0.9461 0.9146 0.6251 0.6551 0.6687
@@ F.1 For RoBERTa
As mentioned in the main text, when measuring the Spearman correlation between the number of intruder dimensions and forgetting in Fig. 7 we find an extremely strong fit, with $\rho=0.971$ and p-value $\ll 0.001$. This shows us that intruder dimensions strongly correlate with forgetting. In contrast, when we correlate intruder dimensions with performance, we find no such correlation: for RoBERTa, we measure $\rho=-0.3381$ with p-value $=0.218$.
@@ F.2 LLaMA2-7B
When measuring the Spearman correlation between number of intruder dimensions and forgetting for our LLaMA2-7B models across training epochs (Fig. 11 (middle)), we find a statistically significant relationship with $\rho=0.59$ and p-value $=0.0006$. When correlating intruder dimensions and test accuracy (Fig. 11 (left)), we instead measure $\rho=-0.3178$ with p-value $=0.0869$. Again, we see that intruder dimensions correlates with forgetting.

### v3

@@ 3 Structural Differences
1:Pre‑trained weights $W_{0}$, fine‑tuned weights $W_{\text{t}}$, cosine similarity threshold $\epsilon$, # of fine-tuned singular vectors to examine $k$.
[…]
@@
LoRA fine-tuned models contain high-ranking intruder dimensions while fully fine-tuned models do not. To characterize the differences in fine-tuning methods, we first evaluate the differences in the total number of intruder dimensions in the top 10 highest-ranking singular vectors ($k=10$). We repeat this procedure for a range of $\epsilon$ values, our cosine similarity threshold. The results are presented in Fig. 4. For LLaMA2-7B, we find that models trained with LoRA contain intruder dimensions for ranks at least as high as $r\leq 256$. For RoBERTa, we consistently observe intruder dimensions for rank $r\leq 16$, even for low values of $\epsilon$. Interestingly, we observe that fully fine-tuned models, for all model sizes, almost never contain intruder dimensions in their top 10 singular vectors, even for epsilon values of about 0.6 to 0.9. This means that full fine-tuning makes smaller changes to the same set of high contribution pre-trained singular vectors, rather than introducing new singular vectors like LoRA. Importantly, the number of intruder dimensions appears to drop as rank increases past a certain threshold, suggesting that the low-rank nature, as well as the update rule of LoRA, induces them to occur. This is underscored by the $r=2048$ case of LLaMA2-7B fine-tuned on math (Fig 4(b)), which does not have intruder dimensions and instead has a very similar curve to full fine-tuning. As rank increases past a threshold and LoRA begins to resemble a high rank update, intruder dimensions begin to disappear.
LoRA variants have intruder dimensions. We examine 4 other LoRA variants to ensure that our findings do not only apply to vanilla LoRA. We examine AdaLoRA (Zhang et al., 2023), LoRA+ (Hayou et al., 2024), PiSSA (Meng et al., 2024), and VeRA (Kopiczko et al., 2024). In all of these cases, we find intruder dimensions with similar characteristics to vanilla LoRA (see Fig. 23). This shows our findings hold to other variants. For more discussion about these methods, see Appendix O.
@@ 4 Model Differences: Forgetting and Out-of-Distribution Generalization
LoRA forgets less. We measure the change in out of distribution performance (forgetting) induced by fine-tuning. For LLaMA2-7B, we follow Biderman et al. (2024) and measure forgetting as the average score on Hellaswag (Zellers et al., 2019), WinoGrande (Sakaguchi et al., 2021), Arc-Challenge (Clark et al., 2018). For RoBERTa-base, we measure its “pseudo-loss”, which is analogous to language modelling loss for encoder-only models, as described by Salazar et al. (2020) on a sample of its pre-training dataset (as described by Liu et al. (2019)). Going forward, we refer to these values as “forgetting” and report them in Fig. 6. We observe that across all tasks, full fine-tuning forgets more of its pre-training language modeling ability in comparison to LoRA. Importantly, all our RoBERTa-base models fine-tune to equivalent accuracy on the downstream task (Table 2). This extends the finding that LoRA forgets less (Biderman et al., 2024) to the case where LoRA and full fine-tuning have equal fit, showing that LoRA forgetting less is not simply a function of it underfitting the fine-tuning task in comparison to full fine-tuning (like in Biderman et al. (2024)), but rather a characteristic of LoRA itself.
@@
Intruder dimensions strongly correlate with forgetting. When we measure the Spearman correlation between the number of intruder dimensions with forgetting in Fig. 7, we find an extremely strong fit($\rho=0.971$, p-value $\ll 0.001$). When measuring the same for our LLaMA2-7B models across training epochs (Fig. 11), we still find a strong and still statistically significant relationship ($\rho=0.59$, p-value =0.0006=0.0006). In contrast, when measuring the correlation between intruder dimensions and test accuracy, we find no statistically significant relationship: for RoBERTa, we measure $\rho=-0.3381$ & p-value =0.218=0.218. For LLaMA2-7B, we measure $\rho=-0.3178$ & p-value =0.0869=0.0869. See Appendix F for more information. These results suggest that intruder dimensions are clearly linked with forgetting but are not necessary for performance. We examine this claim and whether this relationship is causal in the next section.
@@ A.2 Adding a random vector
[…]
@@ Appendix C Model Accuracies
RoBERTa-base Full 0.8745 0.9438 0.9152 0.6582 0.6499 0.6892
@@ F.1 For RoBERTa
As mentioned in the main text, when measuring the Spearman correlation between the number of intruder dimensions and forgetting in Fig. 7 we find an extremely strong fit, with $\rho=0.971$ and p-value $\ll 0.001$. This shows us that intruder dimensions strongly correlate with forgetting. In contrast, when we correlate intruder dimensions with performance, we find no such correlation: for RoBERTa, we measure $\rho=-0.3381$ with p-value =0.218=0.218.
@@ F.2 LLaMA2-7B
When measuring the Spearman correlation between number of intruder dimensions and forgetting for our LLaMA2-7B models across training epochs (Fig. 11 (middle)), we find a statistically significant relationship with $\rho=0.59$ and p-value =0.0006=0.0006. When correlating intruder dimensions and test accuracy (Fig. 11 (left)), we instead measure $\rho=-0.3178$ with p-value =0.0869=0.0869. Again, we see that intruder dimensions correlates with forgetting.